OpenStudy (anonymous):

The ratio of two numbers is 3:b 7, if 4 is added to the greater number and 4 is subtracted from the smaller number , the greater number becomes three times the smaller one.find the numbers. A) 12,28 B) B18,42 C) 24,56 D) 21,49 E) 15,35

OpenStudy (anonymous):

Ratio of two number is what???

OpenStudy (anonymous):

Is it 3:7??

OpenStudy (anonymous):

Let the one number be: 3x So, other will be : 7x 7x + 4 = 3(3x - 4) 7x + 4 = 9x - 12 2x = 16 Find x from here.. And then find the numbers that are 3x and 7x..

OpenStudy (anonymous):

sorry that's 3: 7 typo

OpenStudy (anonymous):

2x = 16 Divide by 2 both the sides and get x.. Then find 3x and 7x..